Hydrogen gas can be prepared by reaction of zinc metal with aqueous HCl: Zn(s)+2HCl(aq)→ZnCl2(aq)+H2(g)

How many liters of H2 would be formed at 680 mm Hg and 17 ∘C if 27.0 g of zinc was allowed to react?

How many grams of zinc would you start with if you wanted to prepare 6.00 L of H2 at 400 mm Hg and 35.0 ∘C?
Answer #1

From the equation of the reaction, we see that the number of moles of H2 formed is equal to the number of moles of Zn that reacts. Therefore, we find the number of moles of Zn.

Number of moles = mass / molar mass = 27 / 65.98 = 0.41 moles

Number of moles of H2 = 0.41 moles

From the ideal gas equation, we have: V = n * R * T / P

Using the values in the question, converting Pressure to atm and temperature to Kelvin, we have:

V = 0.41 * 0.0821 * 290 / [680/760]

V = 10.91 L

The next part of the problem is simply the reverse of what we did here, i.e finding the mass of Zn from Volume of H2. We first find the number of moles of H2

n = P * V / [R *T]

n = (400/760) * 6 / [0.0821 * 308]

n = 0.125 moles

This is equal to the number of moles of Zn that reacts.

Mass of Zn = number of moles * molar mass = 0.125 * 65.98 = 8.25 g